OpenStudy (anonymous):

When there is a percentage of %0.90 and you subtract %0.27. How do you make it into a fraction? @sunnyshores

OpenStudy (shamil98):

0.9 = 9/10 or 90/100 0.27 = 27/100

OpenStudy (anonymous):

she asked me -.- ugh

OpenStudy (shamil98):

\[\huge \frac{ 90 }{ 100 } - \frac{ 27 }{ 100 } = \frac{ 63 }{ 100 }\]

OpenStudy (shamil98):

you're a noob, hush.

OpenStudy (anonymous):

and i needed that medal too....

OpenStudy (anonymous):

wtf im not a noob at math

OpenStudy (anonymous):

lol @shamil98

OpenStudy (shamil98):

\[\huge \frac{ 63 }{ 100 } = 0.63 \text{%}\]

OpenStudy (anonymous):

u guys shud try being nice...

OpenStudy (charlotte123):

@shamil98 \(\huge\color{black}{\frak{Noob}}\)

OpenStudy (anonymous):

^ lol

OpenStudy (shamil98):

It's okay, she's an "acquaintance".

OpenStudy (shamil98):

Compared to me, Jenna, you're a noob :)

OpenStudy (anonymous):

b**** -.-